OpenStudy (anonymous):

Can anyone help me through steps how to solve this equation? (Polynomial) 8x^3-1=0

OpenStudy (anonymous):

8x^3-1=0 +1 +1 8x^3=1 /8 /8 x^3=1/8 cubic root on both sides x=cubic root of 1/8

OpenStudy (anonymous):



OpenStudy (anonymous):

x=1/2